Richard Gaziano is a manager for Health Care, Inc. Health Care deducts Social Security, Medicare, and FIT (by percentage method) from his earnings. Assume a rate of 6.2% on $128,400 for Social Security and 1.45% for Medicare. Before this payroll, Richard is $1,000 below the maximum level for Social Security earnings. Richard is married, is paid weekly, and claims 2 exemptions.


What is Richard’s net pay for the week if he earns $1,900?
Answer #1

Answer: $1577

Calculation:

(a). Gross Weekly pay before deductions = 1900

(b). Social security tax on $1000 of gross pay = $1000 x .062 = $62

(c). Medicare tax on the gross pay = $1900*.0145 = $27.55

(d). Withholdings from his weekly pay :

Withholding amount for someone who is paid weekly with one exemption is $76.9 . Richard has 2 exemptions, so we multiply by 2: $76.9 x 2 = $153.8, $1900- $153.8 = 1746.2

Tax withholding for a married person who is paid weekly = married person paid weekly (Table 2) who receives Over $1606 but not over $3073 is taxed $198.40 plus 25% of any excess over $1606.

So, 1746.2- 1606 = 140.2

He will pay the tax of $198.4 and the 25% tax on the amount calculated as 140.2,

=198.4+(25%*140.2) = 233.45

e. Deduct (b),(c), and (d) from gross pay(a)

that is $1900 (gross pay) - $62(social security tax) - $27.55(med tax) - $233.45 (federal tax withholding) = $1577

So Net pay is $1577